# Generating Text from Uniform Meaning Representation

### v2

@@ 1 Introduction
Figure 1: UMR graph for the sentence “He was searching for a clue” in graph form (top) and in ‘PENMAN’ notation (kasper-1989-flexible) (bottom).
Uniform Meaning Representation (UMR) is a graph-based semantic representation designed to capture the core elements of meaning for a wide range of languages, accounting for some languages’ limited resources and linguistic diversity in the annotation process (van2021designing). An example sentence-level UMR graph can be seen in Figure˜1. The UMR project has seen recent progress but is still in the early stages, with new resources such as an online web annotation tool (ge-etal-2023-umr) and a relatively small recently-released annotated dataset (bonn-etal-2024-building-broad).
Abstract Meaning Representation (AMR; banarescu-etal-2013-abstract), on which UMR is based, is an annotation schema designed for English which has seen success and adoption by the broader NLP community, including incorporation into downstream applications and studies of language (wein-opitz-2024-survey); this is in large part due to the substantial efforts made towards high-quality text-to-AMR parsing and AMR-to-text generation models (sadeddine-etal-2024-survey). [^2] The breadth of parsing and generation work for AMR can be seen in the AMR Bibliography. Thus, in order to see similar success for the multilingual UMR schema, efforts towards UMR-to-text generation and text-to-UMR parsing are critical. As such, the multilingual (and document-level) benefits of UMR cannot be exploited by the NLP community until the tools exist to easily parse and generate.
@@ 2 Uniform Meaning Representation
UMR is based on Abstract Meaning Representation (AMR), which was designed for English (banarescu-etal-2013-abstract) but has since seen various individual language extensions and cross-lingual applications (10.1162/coli_a_00503). While AMR requires these individual language-specific adaptations to accommodate the linguistic features of the individual language (such as pronoun drop, grammatical number, and affixes), UMR’s flexibility enables annotation of many languages within the same schema, showing its promise as a multilingual representation (wein-bonn-2023-comparing).
@@
Similarly to AMR, UMR annotations are rooted, directed graphs that capture meaning (van2021designing). Unlike AMR, UMR additionally incorporates aspect and modality at the sentence-level, contains document-level information (enabling annotation of coreferential relations), and alignments between the coreferential elements.
@@
In order to ensure that UMR could reflect meaning for many languages, the UMR schema is flexible in its annotation while ensuring consistency across languages, differentiating it from the English-based AMR. At the sentence-level, UMR accounts for linguistic diversity across languages through the use of a lattice-like annotation schema (van-gysel-etal-2019-cross), enabling annotators to apply a more coarse-grained or fine-grained annotation based on the features of the language being annotated. Particular care is also given to the annotation of low- or no-resource languages (vigus-etal-2020-cross), as “Stage 0” annotation is performed for languages that do not have existing rolesets available. Stage 0 annotation enables annotators to establish predicate-argument structures and develop rolesets while performing UMR annotation.
@@
While text-to-AMR parsers and AMR-to-text generation models have seen substantial progress in recent years (with Smatch scores of 86% for parsing (lee-etal-2022-maximum; vasylenko-etal-2023-incorporating) and 50 BLEU points for generation (cheng-etal-2022-bibl)), no efforts have yet been made towards UMR-to-text generation models, largely due to the recent development of the schema. We opt to use AMR-to-text models as a baseline for UMR-to-text generation, though the fact that UMR contains more semantic information than AMR (such as aspect, modality, and document-level information) makes AMR-to-text models not particularly well-suited for this task. Models trained solely on AMR are unlikely to correctly account for these elements, leading to a loss of critical semantic information. The use of AMR-to-text generation models as a baseline motivates progress on dedicated UMR-to-text generation approaches, such as ours.
@@
Recent work has investigated automated Arápaho UMR annotation (buchholz-etal-2024-bootstrapping-umr), automatic annotation of tense and aspect for UMR (chen-etal-2021-autoaspect), and conversion of AMR graphs to UMR (bonn-etal-2023-mapping; post-etal-2024-accelerating). Special attention has been given to UMR annotation of multi-word expressions (bonn-etal-2023-umr) and Chinese verb compounds (sun-etal-2023-umr).
@@
Contributing to the UMR technological ecosystem, chun-xue-2024-pipeline released a text-to-UMR parser, which produces the document-level UMR graph based on the contents of the sentence, and the sentence-level UMR graph by running existing text-to-AMR parsers and then converting the AMR into a UMR. This pipeline parsing approach mirrors our baseline approach, in the reverse direction.
@@ 3.1 Data
We use the first release of UMR data (bonn-etal-2024-building-broad), which contains annotations in English, Chinese, and four languages indigenous to the Americas: Arápaho, Navajo, Sanapaná, and Kukama. The English data contains LORELEI news text and a description of a silent film. The Chinese data consists of wikinews sentences. The Arápaho, Navajo, and Kukama annotations all represent narrative documents, while it is not clear what genre the Sanapaná annotations are (bonn-etal-2024-building-broad). We first take out the English UMR data that contains equivalent AMRs in the AMR3.0 dataset (knight_amr_3.0), as to avoid leakage and unfair evaluations of models which contain the corresponding AMR data, removing 66 English UMRs.
@@ 3.2 Automatic Evaluation
We compare the generated text from each of our approaches against the reference sentences by using BERTscore (zhang2019bertscore), given its previously evidenced correlation with human judgments for AMR-to-English text generation (manning-etal-2020-human). We specifically use multilingual BERTscore (mBERTscore), as well as METEOR (banerjee-lavie-2005-meteor) and BLEU (papineni2002bleu) to enable ease of comparison with future work.
@@ 3.4 Human Evaluation
We evaluated the inter-annotator agreement using the Pearson correlation coefficient (Pearson1895), calculating pairwise agreement for each of the four surveys (Chinese and English, fluency and adequacy for each). The average correlation coefficient for English fluency was 0.72, for English adequacy was 0.78, for Chinese fluency was 0.55, and for Chinese adequacy was 0.64. This indicates that there is a strong correlation between the annotators’ scores, validating the judgments, while the English raters exhibited even greater agreement. [^3] This is likely due to the generated text being both more fluent and more adequate for English than Chinese (see Sections 5.2 and 6.2). Note also that all pair-wise correlation coefficients had a statistically significant p-value ($p<0.05$).
@@ 4.1 Models
    amrlib555amrlib GitHub Repository: sequence-to-sequence T5 model, trained on AMR3.0 (knight_amr_3.0)
@@
    AMRBART (bai-etal-2022-graph): BART-based (lewis-etal-2020-bart) model pretrained on English text, AMR graphs from AMR2.0 (knight_amr_2.0) and AMR3.0, along with 200k silver AMRs parsed by SPRING
@@
    SPRING2 and SPRING3 (bevilacqua-etal-2021-one): BART-based sequence-to-sequence model trained on AMR v3.0 that simplifies AMR parsing and generation by treating them as symmetric tasks. SPRING2 and SPRING3 differ in their training datasets: SPRING2 is trained on AMR2.0 while SPRING3 is trained on AMR3.0
@@
    BiBL (cheng-etal-2022-bibl): utilizes the architectural framework of SPRING to align AMR graphs and text, in order to share information across the parsing and generation tasks trained on AMR2.0 and AMR3.0
@@
    Smelting (ribeiro-etal-2021-smelting): trained mT5 (xue2021mt5) on gold English AMR graphs and sentences as well as silver machine-translated sentences
@@ 5.1 Conversion Process
In order to convert UMRs into AMRs, we design a rule-based conversion process. During conversion, some elements of UMR do not appear in AMR (such as aspect and mode) and are thus simply removed. Other changes include converting split roles, renaming roles, and introducing additional roles. This process is informed by bonn-etal-2023-mapping as well as the AMR (AMRGuidelines) and UMR guidelines (UMRGuidelines). The “person” concepts (:refer-number or refer-person) and pronouns generally are handled in a more complicated way in UMR than AMR, so in order to convert these concepts into their AMR counterparts, we changed all “person” concepts to a single node of the equivalent English pronoun. We convert third-person pronouns, both singular and plural, to “they,” opting for a gender-neutral approach. Additionally, if a graph refers to the first person but did not include a refer-number tag, we default to using the singular pronoun “I.” The Sanapaná data also includes a number of instances of :wiki, which we remove. In total, this process results in 4 kinds of UMR roles being removed and over 65 kinds of roles converted to their equivalent AMR role. [^7] We also map :material to :consist-of, :concessive-condition to :condition, and have-91 to have-03.
@@ 6.1 Methods
We determine which AMR-to-text generation models to fine-tune based on the results from Sections˜4 and 5, picking the best-performing models for the baseline and pipeline approaches. Accordingly, the first model we fine-tune is amrlib, which also follows wein-2022-spanish (which fine-tunes the t5wtense AMR-to-English generation model for Spanish text generation). Next, we fine-tune SPRING2 and Smelting, given their quantitative performances for English and Chinese, respectively. Finally, we fine-tune BiBL due to its high scores, relative brevity, and perceived fluency (as discussed in Section˜4.2).
Then, in order to investigate whether the inclusion of AMR data in the fine-tuning process aids model performance, we fine-tune large language models with no prior exposure to AMR (as opposed to fine-tuning the AMR-to-text generation models). [^8] Granted, it is possible that these models have been exposed to AMR graphs in pretraining, though prior work has indicated that foundation models are unable to parse AMR graphs out-of-the-box (ettinger-etal-2023-expert). These models are mT5 (xue2021mt5), mBART (liu2020multilingual), and Gemma 2B (gemma2024improving).
@@ 6.2 Results
Third, the AMR-to-text fine-tuned models generally perform better at a specific language when they are only trained on that language’s data versus being trained on all languages. For example, English generation quality drops slightly when trained on multilingual corpora compared to English-only data. This may be due to the “curse of multilinguality:” exposure to additional languages can degrade monolingual abilities (conneau-etal-2020-unsupervised). Chinese models in particular exhibit more variability and, on average, perform slightly below their English counterparts across automatic and human metrics. This may reflect lower-quality annotations or differences in UMR complexity.

### v3

@@ 1 Introduction
Figure 1: UMR graph for the sentence “He was searching for a clue” in graph form (top) and in ‘PENMAN’ notation (Kasper, 1989) (bottom).
Uniform Meaning Representation (UMR) is a graph-based semantic representation designed to capture the core elements of meaning for a wide range of languages, accounting for some languages’ limited resources and linguistic diversity in the annotation process (Van Gysel et al., 2021). An example sentence-level UMR graph can be seen in Figure˜1. The UMR project has seen recent progress but is still in the early stages, with new resources such as an online web annotation tool (Ge et al., 2023) and a relatively small recently-released annotated dataset (Bonn et al., 2024).
Abstract Meaning Representation (AMR; Banarescu et al., 2013), on which UMR is based, is an annotation schema designed for English which has seen success and adoption by the broader NLP community, including incorporation into downstream applications and studies of language (Wein and Opitz, 2024); this is in large part due to the substantial efforts made towards high-quality text-to-AMR parsing and AMR-to-text generation models (Sadeddine et al., 2024). [^2] The breadth of parsing and generation work for AMR can be seen in the AMR Bibliography. Thus, in order to see similar success for the multilingual UMR schema, efforts towards UMR-to-text generation and text-to-UMR parsing are critical. As such, the multilingual (and document-level) benefits of UMR cannot be exploited by the NLP community until the tools exist to easily parse and generate.
@@ 2 Uniform Meaning Representation
UMR is based on Abstract Meaning Representation (AMR), which was designed for English (Banarescu et al., 2013) but has since seen various individual language extensions and cross-lingual applications (Wein and Schneider, 2024). While AMR requires these individual language-specific adaptations to accommodate the linguistic features of the individual language (such as pronoun drop, grammatical number, and affixes), UMR’s flexibility enables annotation of many languages within the same schema, showing its promise as a multilingual representation (Wein and Bonn, 2023).
@@
Similarly to AMR, UMR annotations are rooted, directed graphs that capture meaning (Van Gysel et al., 2021). Unlike AMR, UMR additionally incorporates aspect and modality at the sentence-level, contains document-level information (enabling annotation of coreferential relations), and alignments between the coreferential elements.
@@
In order to ensure that UMR could reflect meaning for many languages, the UMR schema is flexible in its annotation while ensuring consistency across languages, differentiating it from the English-based AMR. At the sentence-level, UMR accounts for linguistic diversity across languages through the use of a lattice-like annotation schema (Van Gysel et al., 2019), enabling annotators to apply a more coarse-grained or fine-grained annotation based on the features of the language being annotated. Particular care is also given to the annotation of low- or no-resource languages (Vigus et al., 2020), as “Stage 0” annotation is performed for languages that do not have existing rolesets available. Stage 0 annotation enables annotators to establish predicate-argument structures and develop rolesets while performing UMR annotation.
@@
While text-to-AMR parsers and AMR-to-text generation models have seen substantial progress in recent years (with Smatch scores of 86% for parsing (Lee et al., 2022; Vasylenko et al., 2023) and 50 BLEU points for generation (Cheng et al., 2022)), no efforts have yet been made towards UMR-to-text generation models, largely due to the recent development of the schema. We opt to use AMR-to-text models as a baseline for UMR-to-text generation, though the fact that UMR contains more semantic information than AMR (such as aspect, modality, and document-level information) makes AMR-to-text models not particularly well-suited for this task. Models trained solely on AMR are unlikely to correctly account for these elements, leading to a loss of critical semantic information. The use of AMR-to-text generation models as a baseline motivates progress on dedicated UMR-to-text generation approaches, such as ours.
@@
Recent work has investigated automated Arápaho UMR annotation (Buchholz et al., 2024), automatic annotation of tense and aspect for UMR (Chen et al., 2021), and conversion of AMR graphs to UMR (Bonn et al., 2023b; Post et al., 2024). Special attention has been given to UMR annotation of multi-word expressions (Bonn et al., 2023a) and Chinese verb compounds (Sun et al., 2023).
@@
Contributing to the UMR technological ecosystem, Chun and Xue (2024) released a text-to-UMR parser, which produces the document-level UMR graph based on the contents of the sentence, and the sentence-level UMR graph by running existing text-to-AMR parsers and then converting the AMR into a UMR. This pipeline parsing approach mirrors our baseline approach, in the reverse direction.
@@ 3.1 Data
We use the first release of UMR data (Bonn et al., 2024), which contains annotations in English, Chinese, and four languages indigenous to the Americas: Arápaho, Navajo, Sanapaná, and Kukama. The English data contains LORELEI news text and a description of a silent film. The Chinese data consists of wikinews sentences. The Arápaho, Navajo, and Kukama annotations all represent narrative documents, while it is not clear what genre the Sanapaná annotations are (Bonn et al., 2024). We first take out the English UMR data that contains equivalent AMRs in the AMR3.0 dataset (Knight et al., 2020), as to avoid leakage and unfair evaluations of models which contain the corresponding AMR data, removing 66 English UMRs.
@@ 3.2 Automatic Evaluation
We compare the generated text from each of our approaches against the reference sentences by using BERTscore (Zhang et al., 2020), given its previously evidenced correlation with human judgments for AMR-to-English text generation (Manning et al., 2020). We specifically use multilingual BERTscore (mBERTscore), as well as METEOR (Banerjee and Lavie, 2005) and BLEU (Papineni et al., 2002) to enable ease of comparison with future work.
@@ 3.4 Human Evaluation
We evaluated the inter-annotator agreement using the Pearson correlation coefficient (Pearson and Galton, 1895), calculating pairwise agreement for each of the four surveys (Chinese and English, fluency and adequacy for each). The average correlation coefficient for English fluency was 0.72, for English adequacy was 0.78, for Chinese fluency was 0.55, and for Chinese adequacy was 0.64. This indicates that there is a strong correlation between the annotators’ scores, validating the judgments, while the English raters exhibited even greater agreement. [^3] This is likely due to the generated text being both more fluent and more adequate for English than Chinese (see Sections 5.2 and 6.2). Note also that all pair-wise correlation coefficients had a statistically significant p-value ($p<0.05$).
@@ 4.1 Models
    amrlib555amrlib GitHub Repository: sequence-to-sequence T5 model, trained on AMR3.0 (Knight et al., 2020)
@@
    AMRBART (Bai et al., 2022): BART-based (Lewis et al., 2020) model pretrained on English text, AMR graphs from AMR2.0 (Knight et al., 2017) and AMR3.0, along with 200k silver AMRs parsed by SPRING
@@
    SPRING2 and SPRING3 (Bevilacqua et al., 2021): BART-based sequence-to-sequence model trained on AMR v3.0 that simplifies AMR parsing and generation by treating them as symmetric tasks. SPRING2 and SPRING3 differ in their training datasets: SPRING2 is trained on AMR2.0 while SPRING3 is trained on AMR3.0
@@
    BiBL (Cheng et al., 2022): utilizes the architectural framework of SPRING to align AMR graphs and text, in order to share information across the parsing and generation tasks trained on AMR2.0 and AMR3.0
@@
    Smelting (Ribeiro et al., 2021): trained mT5 (Xue et al., 2021) on gold English AMR graphs and sentences as well as silver machine-translated sentences
@@ 5.1 Conversion Process
In order to convert UMRs into AMRs, we design a rule-based conversion process. During conversion, some elements of UMR do not appear in AMR (such as aspect and mode) and are thus simply removed. Other changes include converting split roles, renaming roles, and introducing additional roles. This process is informed by Bonn et al. (2023b) as well as the AMR (Banarescu et al., 2019) and UMR guidelines (30). The “person” concepts (:refer-number or refer-person) and pronouns generally are handled in a more complicated way in UMR than AMR, so in order to convert these concepts into their AMR counterparts, we changed all “person” concepts to a single node of the equivalent English pronoun. We convert third-person pronouns, both singular and plural, to “they,” opting for a gender-neutral approach. Additionally, if a graph refers to the first person but did not include a refer-number tag, we default to using the singular pronoun “I.” The Sanapaná data also includes a number of instances of :wiki, which we remove. In total, this process results in 4 kinds of UMR roles being removed and over 65 kinds of roles converted to their equivalent AMR role. [^7] We also map :material to :consist-of, :concessive-condition to :condition, and have-91 to have-03.
@@ 6.1 Methods
We determine which AMR-to-text generation models to fine-tune based on the results from Sections˜4 and 5, picking the best-performing models for the baseline and pipeline approaches. Accordingly, the first model we fine-tune is amrlib, which also follows Wein et al. (2022) (which fine-tunes the t5wtense AMR-to-English generation model for Spanish text generation). Next, we fine-tune SPRING2 and Smelting, given their quantitative performances for English and Chinese, respectively. Finally, we fine-tune BiBL due to its high scores, relative brevity, and perceived fluency (as discussed in Section˜4.2).
Then, in order to investigate whether the inclusion of AMR data in the fine-tuning process aids model performance, we fine-tune large language models with no prior exposure to AMR (as opposed to fine-tuning the AMR-to-text generation models). [^8] Granted, it is possible that these models have been exposed to AMR graphs in pretraining, though prior work has indicated that foundation models are unable to parse AMR graphs out-of-the-box (Ettinger et al., 2023). These models are mT5 (Xue et al., 2021), mBART (Liu et al., 2020), and Gemma 2B (Gemma Team, 2024).
@@ 6.2 Results
Third, the AMR-to-text fine-tuned models generally perform better at a specific language when they are only trained on that language’s data versus being trained on all languages. For example, English generation quality drops slightly when trained on multilingual corpora compared to English-only data. This may be due to the “curse of multilinguality:” exposure to additional languages can degrade monolingual abilities (Conneau et al., 2020). Chinese models in particular exhibit more variability and, on average, perform slightly below their English counterparts across automatic and human metrics. This may reflect lower-quality annotations or differences in UMR complexity.